The position of a particle moving in a straight line is give by X=3+4t+3².where x in metre and time in second.Find the values of the physical quantities of the particle at t=2s
1.Position
2.Displacement
3.velocity
4.Acceleration
1
Expert's answer
2018-01-22T14:06:08-0500
At the t=2 s
Position x(2)=3+4×2+3×2^2=23 m
Displacement s= x(2)-x(0)=23-3=20 m
Velocity v=x^' (2)=├ 4+6t┤|_(t=2)=4+6×2=16 m/s
Acceleration a=v^' (t)=6 m/s^2